# An uncountable subring of $\mathbb{R}$ with Hausdorff dimension zero

### v1

@@ Abstract.
We construct an uncountable subring of $\mathbb{R}$ with Hausdorff dimension zero (and hence of Lebesgue measure zero).
We denote by $\mathbb{N}_{0}$ the set of non-negative integers, by $\mathbb{N}$ the set of positive integers and by $\mathbb{Z}$ the set of integers. Let
$S:=\{0\}\cup\{2^{n}:n\in\mathbb{N}_{0}\}=\{0,1,2,2^{2},2^{3},\ldots\}.$
For every $n\in\mathbb{N}$ by $nS$, we denote the $n^{\text{th}}$ sumset of $S$, i.e.,
@@
$A_{n}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}:(x_{k})_{k\in nS}\text{is a bounded sequence of integers}\right\}.$
@@
We first show that $A_{n}$ is a subgroup of $\mathbb{R}$. Trivially, $0\in A_{n}$ and $-x\in A_{n}$ whenever $x\in A_{n}$. Thus, it suffices to show that $A_{n}$ is closed under addition. Let $x,y\in A_{n}$. Therefore,
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{k\in nS}\frac{y_{k}}{2^{k}}$
where $(x_{k})_{k\in nS}$ and $(y_{k})_{k\in nS}$ are bounded sequences of integers. It follows that
@@
is also an element of $A_{n}$.
Clearly, the set $A_{n}$ is uncountable, because by the uniqueness of binary expansions which do not end with an infinite string of $1$’s, the set of real numbers of the form
$\sum\limits_{k\in nS}\frac{z_{k}}{2^{k}}\quad\mbox{with }z_{k}\in\{0,1\}$
@@
$A_{n,t}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\in\mathbb{R}:x_{k}\in\mathbb{Z}\mbox{ and }|x_{k}|\leq t\mbox{ for all }k\in nS\right\}.$
We will need the following result to show that the set $A_{n,t}$ has Hausdorff dimension zero for every $n,t\in\mathbb{N}$.
@@
For $n\in\mathbb{N}$ and $z\geq 1$, define
$g_{n}(z):=\emph{{\#}}\{k\in nS:k\leq z\}.$
@@
Clearly, for every $z\geq 1$, we have
$g_{n}(z)\leq g_{1}(z)^{n}\mbox{ and }g_{1}(z)\leq 2+\text{log}_{2}z.$
@@
Write $nS=\{k_{m}:m\in\mathbb{N}\}$ where $(k_{m})_{m\in\mathbb{N}}$ is an increasing sequence. Let $x\in nS$ and write
$x=\sum\limits_{i=1}^{\infty}\frac{a_{i}}{2^{k_{i}}}$
with $a_{i}\in\mathbb{Z}$ and $|a_{i}|\leq t$. Pick $j\in\mathbb{N}$ and split $x$ into
$x=m_{x}+e_{x},$
@@
$m_{x}:=\sum\limits_{i=1}^{j-1}\frac{a_{i}}{2^{k_{i}}}\quad\mbox{and}\quad e_{x}:=\sum\limits_{i=j}^{\infty}\frac{a_{i}}{2^{k_{i}}}.$
@@
$|e_{x}|\leq t\sum\limits_{i=j}^{\infty}\frac{1}{2^{k_{i}}}\leq t\sum\limits_{h=k_{j}}^{\infty}\frac{1}{2^{h}}=\frac{t}{2^{k_{j}-1}}.$
Let $r_{j}=t/2^{k_{j}-1}$. It follows that
$A_{n,t}\subseteq\bigcup_{m\in X}I_{m},$
where
$X:=\left\{\sum\limits_{i=1}^{j-1}\frac{a_{i}}{2^{k_{i}}}:a_{i}\in\mathbb{Z},\ |a_{i}|\leq t\right\}$
and
$I_{m}:=[m-r_{j},m+r_{j}].$
Then #$\emph{{\#}}(X)\leq(2t+1)^{j-1}$ and for every $m\in X$ we see that $\mu(I_{m})=2r_{j}$ where $\mu(I_{m})$ is the length of the interval $I_{m}$.
Let $d>0$ be a real number, let $\delta_{j}:=4r_{j}$, and let $\mathcal{H}_{\delta_{j}}^{d}(A_{n,t})$ be defined as in [2]. We see that
$\mathcal{H}_{\delta_{j}}^{d}(A_{n,t})\leq\sum\limits_{m\in X}\mu(I_{m})^{d}=\emph{{\#}}(X)(2r_{j})^{d}\leq\frac{(2t+1)^{j-1}t^{d}}{2^{(k_{j}-2)d}}.$
@@
(1) $\lim\limits_{j\rightarrow\infty}\frac{(2t+1)^{j-1}t^{d}}{2^{(k_{j}-2)d}}=0.$
Since $\delta_{j}\to 0$ as $j\to\infty$, this implies that $\mathcal{H}^{d}(A_{n,t})=0$ where $\mathcal{H}^{d}(A_{n,t})$ is the $d$-dimensional Hausdorff measure of $A_{n,t}$. (For the definition of this measure and related terms, see [2].) Since this holds for all $d>0$, we conclude that the Hausdorff dimension of $A_{n,t}$ is zero.
@@
$\lim\limits_{j\rightarrow\infty}\frac{(3t)^{j}}{2^{k_{j}d}}=0.$
By taking logarithms, we see that (1) holds if
$\lim\limits_{j\rightarrow\infty}\big{(}k_{j}d-j\log_{2}(3t)\big{)}=\infty.$
As $d>0$, we let $c$ be the constant $(1/d)\log_{2}(3t)$ and we see that (1) holds if
(2) $\lim\limits_{j\to\infty}(k_{j}-cj)=\infty.$
@@
$k_{j}\geq 2^{\sqrt[n]{j}-2}.$
Hence, (2) follows, which justifies our claim. ∎
@@
We deduce that the set $A$ is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. We aim to establish that $A$ is in fact a subring. To this end, we need the following result.
@@
Let $n\in\mathbb{N}$ and let $a_{n}(s)$ be the number of representations of an element $s\in\mathbb{N}$ as a sum $s=t_{1}+t_{2}+\cdots+t_{n}$ of $n$ elements $t_{1},t_{2},\ldots,t_{n}\in S$. Then there exists $c_{n}\in\mathbb{R}$ such that $a_{n}(s)\leq c_{n}$ for every $s\in\mathbb{N}$.
@@
Clearly $a_{1}(s)\leq 1$ for every $s\in\mathbb{N}$ and hence, we may take $c_{1}:=1$. Now assume that $n\geq 2$ and that there exists $c_{n-1}\in\mathbb{R}$ such that $a_{n-1}(s)\leq c_{n-1}$ for every $s\in\mathbb{N}$. Let $s\in\mathbb{N}$ and suppose
$s=t_{1}+\cdots+t_{n}$
is a representation of $s$ as a sum of $n$ elements of $S$. Then for at least one $i\in\{1,\ldots,n\}$, we have $t_{i}\geq s/n$ and hence $t_{i}\neq 0$. So there exists $k\in\mathbb{N}_{0}$ such that $t_{i}=2^{k}$. As $s/n\leq t_{i}\leq s$, we observe that $\log_{2}s-\log_{2}n\leq k\leq\log_{2}s$. So if
$B_{n,s}:=[\,\log_{2}s-\log_{2}n,\,\log_{2}s\,]\cap\mathbb{N}_{0},$
@@
$a_{n}(s)\leq n\sum\limits_{k\in B_{n,s}}a_{n-1}(s-2^{k})\leq n(1+\log_{2}n)c_{n-1}.$
So we are done by taking $c_{n}:=n(1+\log_{2}n)c_{n-1}$. ∎
@@
It suffices to show that $1\in A$ and that $A$ is closed under multiplication. We see that
$1=\frac{1}{2^{0}}+\sum\limits_{k\in S\setminus\{0\}}\frac{0}{2^{k}}\in A_{1}.$
To establish that $A$ is closed under multiplication, we proceed as follows. Let $x,y\in A$. There exists $m,n\in\mathbb{N}$ such that $x\in A_{m}$ and $y\in A_{n}$. Therefore,
$x=\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}$
where $(x_{k})_{k\in mS}$ and $(y_{\ell})_{\ell\in nS}$ are bounded sequences of integers. Now
@@
$z_{r}:=\sum\limits_{\begin{subarray}{c}(k,\ell)\in mS\times nS\\ k+\ell=r\end{subarray}}x_{k}y_{\ell}.$
The number of representations of an element $r\in(m+n)S$ as a sum $r=k+\ell$ of elements $k\in mS$ and $\ell\in nS$ is bounded by the number of representations of $r$ which are of the form $r=t_{1}+t_{2}+\cdots+t_{m+n}$ where $t_{1},t_{2},\ldots,t_{m+n}\in S$. Hence, we deduce that the sequence $(z_{r})_{r\in(m+n)S}$ is bounded as a consequence of Lemma 2 and the boundedness of the sequences $(x_{k})_{k\in mS}$ and $(y_{\ell})_{\ell\in nS}$. It follows that $xy\in A_{m+n}$. This establishes that $A$ is closed under multiplication. ∎
We now proceed to show that the only rational numbers in $A$ are the dyadic rationals. In particular, it will follow that the ring $A$ is not a field.
Lemma 3.
Let $r,n\in\mathbb{N}$. Write $nS=\{k_{m}:m\in\mathbb{N}\}$ where $(k_{m})_{m\in\mathbb{N}}$ is an increasing sequence. Then there exists $m\in\mathbb{N}$ such that $k_{m+1}-k_{m}>r$.
Proof.
For every $z\in\mathbb{N}$, let $g_{n}(z):=\emph{{\#}}\{k\in nS:k\leq z\}$. If the conclusion is false, then $g_{n}(z)\geq z/r$ for every $z\in\mathbb{N}$. This contradicts Lemma 1. ∎
Proposition 1.
Let $x\in A$ and $l\in\mathbb{N}$. Then there exists a binary expansion of $x$ which contains either a string of zeroes or a string of ones having length $l$.
Proof.
Let $x\in A_{n,t}$. Then there exists a sequence of integers $(x_{k})_{k\in nS}$ with $|x_{k}|\leq t$ for all $k\in nS$ such that
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}=\sum\limits_{i=1}^{\infty}\frac{x_{k_{i}}}{2^{k_{i}}}.$
Let $m\in\mathbb{N}$. We observe that
$2^{k_{m}}x=q_{m}+r_{m}$
where
$q_{m}:=\sum\limits_{i=0}^{m}2^{k_{m}-k_{i}}x_{k_{i}}\mbox{ and }r_{m}:=\sum\limits_{i=m+1}^{\infty}\frac{x_{k_{i}}}{2^{k_{i}-k_{m}}}.$
@@
$\left|r_{m}\right|\leq~{}t\sum\limits_{j=k_{m+1}-k_{m}}^{\infty}\frac{1}{2^{j}}=\frac{2t}{2^{k_{m+1}-k_{m}}}.$
By Lemma 3, there exists $m\in\mathbb{N}$ such that $2^{k_{m+1}-k_{m}}>2^{l+1}t$ and so $|r_{m}|<1/2^{l}$. Hence,
$\text{frac}\left(2^{k_{m}}x\right)=\begin{cases}r_{m},&\mbox{ if }r_{m}\in\left[\,0,\frac{1}{2^{l}}\right);\\ 1+r_{m},&\mbox{ if }r_{m}\in\left(-\frac{1}{2^{l}},0\right);\end{cases}$
@@
$\text{frac}(2^{k_{m}}x)\in\left[0,\tfrac{1}{2^{l}}\right)\cup\left(1-\tfrac{1}{2^{l}},1\right),$
where frac$(z):=z-[z]$ is the fractional part of $z\in\mathbb{R}$. It follows that the digits after the dot in a binary expansion of $x$ from the $(k_{m}+1)^{\text{th}}$ position to the $(k_{m}+l)^{\text{th}}$ position are all equal. ∎
@@
From Proposition 1, we see that the ring $A$ only contains those rational numbers which have a finite binary expansion. The result follows. ∎
@@
Acknowledgment: We thank Prof. Shrikrishna G. Dani for sharing with us a preprint of [1] in which we see a different construction of an uncountable subgroup of $\mathbb{R}$ having Hausdorff dimension zero.

### v2

@@ Abstract.
We construct a subring as mentioned in the title (hence this subring has Lebesgue measure zero).
1. Preliminaries
For a subset $B\subseteq\mathbb{R}$ we let $d(B)$ denote the diameter of $B$. Given a countable collection of subsets $\mathcal{C}\subseteq\mathcal{P}(\mathbb{R})$ and $s\geq 0$ we define
$d(s,\mathcal{C}):=\sum_{B\in\mathcal{C}}d(B)^{s}.$
We note that $d(s,\mathcal{C})\in[0,\infty]$ and does not depend on the ordering of the subsets in the collection $\mathcal{C}$. For $\delta>0$, we say that a collection of subsets $\mathcal{C}$ is a $\delta$-cover of a subset $F\subseteq\mathbb{R}$ if for every $B\in\mathcal{C}$ we have $d(B)\leq\delta$ and if $F\subseteq\cup\,\mathcal{C}$.
Given a subset $F\subseteq\mathbb{R}$ and $s,\delta\in\mathbb{R}$ with $s\geq 0$, and $\delta>0$, we define
$\mathcal{H}^{s}_{\delta}(F)=\inf\big{\{}d(s,\mathcal{C}):\mathcal{C}\mbox{ is a countable $\delta$-cover of $F$}\big{\}}.$
If $\delta_{1}<\delta_{2}$, then we see that $\mathcal{H}^{s}_{\delta_{1}}(F)\geq\mathcal{H}^{s}_{\delta_{2}}(F)$. The $s$-dimensional Hausdorff measure $\mathcal{H}^{s}(F)$ is defined to be the limit of $\mathcal{H}_{\delta}^{s}(F)$ as $\delta\to 0^{+}$.
It can be shown (see [4]) that there exists $\alpha\in[0,1]$ such that for every $s\in[0,\alpha)$ we have $\mathcal{H}^{s}(F)=\infty$, and for every $s>\alpha$ we have $\mathcal{H}^{s}(F)=0$. The number $\alpha$ is called the Hausdorff dimension of $F$ and is denoted by $\dim_{\mathcal{H}}F$. If $F=\cup\,\mathcal{C}$ where $\mathcal{C}$ is a countable collection of subsets of $\mathbb{R}$, then $\dim_{\mathcal{H}}F=\sup\big{\{}\dim_{\mathcal{H}}A:A\in\mathcal{C}\big{\}}$.
2. Construction
We give an explicit construction of a proper subring of $\mathbb{R}$ which is a Borel set. We also show explicitly that it has Hausdorff dimension zero.
We let $\mathbb{N}_{0}$ denote the set of non-negative integers, $\mathbb{N}$ denote the set of positive integers, and $\mathbb{Z}$ denote the set of integers. For every $a,b\in\mathbb{Z}$ we let $a,b]_{\mathbb{Z}}=\{k\in\mathbb{Z}:a\leq k\leq b\}$. Let
$S:=\{0,1,2,2^{2},2^{3},\ldots\}=\{0\}\cup\{2^{n}:n\in\mathbb{N}_{0}\}.$
For every $n\in\mathbb{N}$ we let the $n^{\text{th}}$ sumset of $S$ be defined as
@@ 2. Construction
[…]
@@
We first show that $A_{n}$ is a subgroup of $\mathbb{R}$. Trivially, we have $0\in A_{n}$, and that $x\in A_{n}$ whenever $x\in A_{n}$. Thus, it suffices to show that $A_{n}$ is closed under addition. Let $x,y\in A_{n}$. Therefore, there exists $t\in\mathbb{N}$ and for every $k\in nS$ there exist $x_{k},y_{k}\in[-t,t\,]_{\mathbb{Z}}$ such that
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{k\in nS}\frac{y_{k}}{2^{k}}.$
It follows that
@@
is also an element of $A_{n}$ since $x_{k}+y_{k}\in[-2t,2t\,]_{\mathbb{Z}}$ for every $k\in nS$. Also, the set $A_{n}$ is uncountable, because by the uniqueness of binary expansions which do not end with an infinite string of ones, the set of real numbers of the form
[…]
@@
$A_{n,t}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\in\mathbb{R}:\mbox{ for every }k\in nS\mbox{ we have }x_{k}\in[-t,t\,]_{\mathbb{Z}}\right\}.$
For every $n,\ell\in\mathbb{N}$ we let $(nS)_{\leq\ell}=\{k\in nS:k\leq\ell\}$ and we let $(nS)_{>\ell}=\{k\in nS:k>\ell\}$. We will need the following result to show that the set $A_{n,t}$ has Hausdorff dimension zero for every $n,t\in\mathbb{N}$.
@@
For every $n,\ell\in\mathbb{N}$ we define
$g_{n}(\ell):=\emph{{\#}}(nS)_{\leq\ell}.$
@@
We observe that for every $z\in\mathbb{N}$, we have
$g_{n}(\ell)\leq g_{1}(\ell)^{n}\mbox{ and }g_{1}(\ell)\leq 2+\log_{2}\ell.$
@@
Let $x\in A_{n,t}$. Then for every $k\in nS$ there exist $a_{k}\in[-t,t\,]_{\mathbb{Z}}$ such that
$x=\sum\limits_{k\in nS}\frac{a_{k}}{2^{k}}.$
Pick $\ell\in\mathbb{N}$ and write $x$ as
$x=y+z,$
@@
$y:=\sum\limits_{k\,\in\,(nS)_{\leq\ell}}\frac{a_{k}}{2^{k}}\quad\mbox{and}\quad z:=\sum\limits_{k\,\in\,(nS)_{>\ell}}\frac{a_{k}}{2^{k}}.$
@@
$|z|\leq t\sum\limits_{k\,\in\,(nS)_{>\ell}}\frac{1}{2^{k}}\leq t\sum\limits_{k=\ell+1}^{\infty}\frac{1}{2^{k}}=\frac{t}{2^{\ell}}.$
Let
$B:=\left\{\sum\limits_{k\,\in\,(nS)_{\leq\ell}}\frac{a_{k}}{2^{k}}:\mbox{ for every }k\in(nS)_{\leq\ell}\mbox{ we have }a_{k}\in[-t,t\,]_{\mathbb{Z}}\right\}$
and let $r=t/2^{\ell}$. It follows that
$A_{n,t}\subseteq\bigcup_{b\in B}I_{b}\mbox{ where }I_{b}=[b-r,b+r]\mbox{ for every }b\in B.$
Let $c_{\ell}=\emph{{\#}}\,(nS)_{\leq\ell}$. Then $\emph{{\#}}B\leq(2t+1)^{c_{\ell}}$ and for every $b\in B$ we see that $\mu(I_{b})=2r$ where $\mu$ is the Lebesgue measure on $\mathbb{R}$. Let $s\in(0,1]$. We see that
$\mathcal{H}_{2r_{\ell}}^{s}(A_{n,t})\leq\sum\limits_{b\in B}\mu(I_{b})^{s}=\emph{{\#}}B\,(2r)^{s}\leq(2t+1)^{c_{\ell}}\frac{t^{s}}{2^{(\ell-1)s}}.$
@@
(1) $\lim\limits_{\ell\rightarrow\infty}(2t+1)^{c_{\ell}}\frac{t^{s}}{2^{(\ell-1)s}}=0.$
Since $r\to 0$ as $\ell\to\infty$, this implies that $\mathcal{H}^{s}(A_{n,t})=0$. As this holds for all $s\in(0,1]$, we conclude that the Hausdorff dimension of $A_{n,t}$ is zero.
@@
$\lim\limits_{\ell\to\infty}\frac{(3t)^{c_{\ell}}}{2^{\ell s}}=0.$
Now, by taking logarithms, we see that (1) holds if
$\lim\limits_{\ell\to\infty}\Big{(}\ell s-c_{\ell}\log_{2}(3t)\Big{)}=\infty.$
Let $\alpha=\frac{\log_{2}(3t)}{s}$. As $s>0$, we see that (1) holds if
(2) $\lim\limits_{\ell\to\infty}\Big{(}\ell-\alpha c_{\ell}\Big{)}=\infty.$
@@
$\ell\geq 2^{c_{\ell}^{r}-2}$
where $r=1/n$. Hence, it follows that (2) holds. This justifies our claim. ∎
@@
We deduce that the set AA is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. We aim to establish that AA is in fact a subring. To this end, we need the following result.
@@
Let $n\in\mathbb{N}$ and let $c_{n}(k)$ be the number of representations of an element $k\in\mathbb{N}$ as a sum $k=s_{1}+s_{2}+\cdots+s_{n}$ of $n$ elements of $S$. Then there exists $b_{n}\in\mathbb{N}$ such that $c_{n}(k)\leq b_{n}$ for every $k\in\mathbb{N}$.
@@
Clearly $c_{1}(k)\leq 1$ for every $k\in\mathbb{N}$ and hence, we may take $b_{1}:=1$. Now assume that $n\geq 2$ and that there exists $b_{n-1}\in\mathbb{N}$ such that $c_{n-1}(k)\leq b_{n-1}$ for every $k\in\mathbb{N}$. Let $k\in\mathbb{N}$ and suppose
$k=s_{1}+\cdots+s_{n}$
is a representation of $k$ as a sum of $n$ elements of $S$. Then for at least one $i\in\{1,\ldots,n\}$ we have $s_{i}\geq k/n$ and hence $s_{i}\neq 0$. So there exists $m\in\mathbb{N}_{0}$ such that $s_{i}=2^{m}$. As $k/n\leq s_{i}\leq k$, we observe that $\log_{2}k-\log_{2}n\leq m\leq\log_{2}k$. Let
$I_{n,k}:=[\,\log_{2}k-\log_{2}n,\,\log_{2}k\,]\cap\mathbb{N}_{0}.$
@@
$c_{n}(k)\leq n\sum\limits_{m\in I_{n,k}}c_{n-1}(k-2^{m})\leq nb_{n-1}\emph{{\#}}I_{n,k}\leq nb_{n-1}(1+\log_{2}n).$
So we are done by taking $b_{n}\in\mathbb{N}$ such that $b_{n}\geq nb_{n-1}(1+\log_{2}n)$. ∎
@@
It suffices to show that $1\in A$ and that AA is closed under multiplication. We see that
$1=\frac{1}{2^{0}}+\sum\limits_{k\in S_{>0}}\frac{0}{2^{k}}\in A_{1}.$
To establish that AA is closed under multiplication, we proceed as follows. Let $x,y\in A$. There exist $m,n\in\mathbb{N}$ such that $x\in A_{m}$ and $y\in A_{n}$. Therefore, there exist $t,t^{\prime}\in\mathbb{N}$ and for every $k\in nS$ there exist $x_{k}\in[-t,t\,]_{\mathbb{Z}}$ and $y_{k}\in[-t^{\prime},t^{\prime}]_{\mathbb{Z}}$ such that
$x=\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}.$
Now
@@
where for every $r\in(m+n)S$ we let
$z_{r}:=\sum\limits_{\begin{subarray}{c}(k,\ell)\in mS\times nS\vskip 3.0pt plus 1.0pt minus 1.0pt\\ k+\ell=r\end{subarray}}x_{k}y_{\ell}.$
The number of representations of an element $r\in(m+n)S$ as a sum $r=k+\ell$ of elements $k\in mS$ and $\ell\in nS$ is bounded by the number of representations of $r$ which are of the form $r=s_{1}+s_{2}+\cdots+s_{m+n}$ where $s_{1},s_{2},\ldots,s_{m+n}\in S$. Hence, by Lemma 2 we deduce that there exists $b\in\mathbb{N}$ such that for every $r\in(m+n)S$ we have $|z_{r}|\leq btt^{\prime}$. It follows that $xy\in A_{m+n}$. This establishes that AA is closed under multiplication. ∎
We now proceed to show that the only rational numbers in AA are the dyadic rationals. In particular, it will follow that the ring AA is not a field.
Lemma 3.
Let $b,n\in\mathbb{N}$. Then there exist two consecutive elements $k,k^{\prime}\in nS$ such that $k^{\prime}-k>b$.
Proof.
For every $\ell\in\mathbb{N}$ we let $c_{\ell}:=\emph{{\#}}\,(nS)_{\leq\ell}$. If the conclusion is false, then $c_{\ell}\geq\ell/b$ for every $\ell\in\mathbb{N}$. This contradicts Lemma 1. ∎
Proposition 1.
Let $x\in A$ and $\ell\in\mathbb{N}$. Then there exists a binary expansion of $x$ which contains either a string of zeroes or a string of ones having length $\ell$.
Proof.
Let $x\in A_{n,t}$. Then for every $k\in nS$ there exists $x_{k}\in[-t,t\,]_{\mathbb{Z}}$ such that
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}.$
Let $m\in nS$. We observe that
$2^{m}x=q_{m}+r_{m}$
@@
$q_{m}:=\sum\limits_{k\in(nS)_{\leq m}}2^{m-k}x_{k}\mbox{ and }r_{m}:=\sum\limits_{k\in(nS)_{>m}}\frac{x_{k}}{2^{k-m}}.$
@@
$|r_{m}|\leq~t\sum\limits_{j=m^{\prime}-m}^{\infty}\frac{1}{2^{j}}=\frac{2t}{2^{m^{\prime}-m}}$
where $m,m^{\prime}\in nS$ are consecutive elements.
By Lemma 3, there exist consecutive elements $k,k^{\prime}\in nS$ such that $2^{k^{\prime}-k}>2^{\ell+1}t$ and so we get $|r_{k}|<\frac{1}{2^{\ell}}$. Hence,
$\text{frac}\left(2^{k}x\right)=\begin{cases}r_{k},&\mbox{ if }r_{k}\in\left[\,0,\frac{1}{2^{\ell}}\right);\\ 1+r_{k},&\mbox{ if }r_{k}\in\left(-\frac{1}{2^{\ell}},0\right);\end{cases}$
@@
$\text{frac}(2^{k}x)\in\left[0,\tfrac{1}{2^{\ell}}\right)\cup\left(1-\tfrac{1}{2^{\ell}},1\right),$
[…]
@@
From Proposition 1, we see that the ring AA only contains those rational numbers which have a finite binary expansion. The result follows. ∎
@@
3. Concluding remarks
In 1966, Erdős and Volkmann [2] showed that for every $d\in[0,1]$ there exists a subgroup of $\mathbb{R}$ which is Borel measurable and has Hausdorff dimension $d$. (Example 12.4 of [4]). In an unpublished work (see [3], pg. 212), using the Continuum Hypothesis, R. O. Davies showed that for every $d\in[0,1]$ there exists a subring of $\mathbb{R}$ which has Hausdorff dimension $d$ and which is not a Borel set. In [1], it is shown that every proper subring of $\mathbb{R}$ which is a Borel set, has Hausdorff dimension zero.
It has come to our attention that our construction has some similarities to that of [2], but our series expansion of real numbers is inspired by the binary expansion. Erdős and Volkmann [2] noted that the question of the Hausdorff dimension of subrings and subfields remained open. We feel that it is of interest to see how Lemma 2 is crucially used to show that we get a subring of $\mathbb{R}$ by using numbers which are represented by special series expansions. Also, we are able to see relatively easily that the set AA has Hausdorff dimension zero.
We have been informed of the following construction. If AA is a compact Cantor set such that for every $n\in\mathbb{N}$ the set $A^{n}$ has Hausdorff dimension zero, then it can be shown that the subring of $\mathbb{R}$ which is generated by AA has Hausdorff dimension zero. However, we are unable to find a reference for the construction of such a set. Our construction is very explicit using series expansions and we use neither the Continuum Hypothesis nor the Axiom of choice.